### 09 Jan 2024 v2

# Search for the population of the low-energy continuum of tetraneutron system was performed for the H(He,Li) and H(He,He)HH+ reactions. Evidence for a hump

Search for the population of the low-energy continuum of tetraneutron system was
performed for the H(He,Li) and
H(He,He)HH+ reactions. Evidence for a hump
in the continuum at about 3.5 MeV was observed in both reactions. The
observed statistics is quite low (6 and up to 40 events) corresponding to very
low cross sections of few microbarns or tens of microbarns. The background
conditions for the H(He,Li) reaction are shown to be good,
favoring the physical nature of the observed events. The
H(He,He)HH+ process transforms to the
H(He,Li reaction in the limit of highest H decay
energies. The population of the low-energy region in the n spectrum is
found to be perfectly correlated with the population of the lowest Li
states in the He+H continuum. Theoretical calculations of He
in a five-body + and of n in a four-body hyperspherical models
are presented. The He wave function is shown to contain strong specific
correlations, which may give rise to very low-energy structures in tetraneutron
continuum in extreme-peripheral reaction scenarios.
Keywords: tetraneutron spectrum; He beam; direct
reactions; cryogenic deuterium target; hyperspherical harmonics; five-body
+ model for He; four-body continuum.
Date: January 9, 2024.
1 Introduction
The search for the the multineutron systems is old, but still unsettled
problem of the low-energy nuclear physics. The first ideas about the possible
multineutron systems stability were expressed in the papers
[1, 2] and since that time reiterated on various
occasions. A detailed account of the multineutron studies history, both
experimental and theoretical was provided in the recent review
[3]. In short, either none or only marginal experimental
evidence, which was never confirmed later, was obtained
for the bound tetraneutron. However, such new attempts continue, and there is a
very recent example [4].
The situation has changed with the recent studies of n population in
reactions with He, where four neutrons can be found in a
spatially-separated neutron-halo configuration. In the
He(He,) reaction
[5] the group with marginal statistics of 4 events was observed
at very low center-of-mass (c.m.) n energy MeV with cross
section of nb. Much more encouraging and intriguing was a very recent
H(He,) experiment [6], where the observation
of the “resonance-like structure” at MeV with MeV
was reported. Around 40 events corresponding to this structure were recorded,
which is statistically quite convincing, and the population cross section is
estimated at the sub-microbarn level.
The issue of bound “multineutron nuclei” was scrutinized in the modern
theoretical approaches Ref. [7, 8, 9]
always with the same result: for the two-body potentials considered fixed, a
radical modification of three-body potential is required, which is not tolerable
as leading to “devastating effect” on the nuclear structure in general. In the
theoretical studies of the n low-energy continuum different evidences were
found for important low-energy effect of n final-state interaction (FSI).
These are: specific trajectories of the -matrix poles, enhanced “time
delay”, energy extrapolations of multi-neutron states confined in traps, etc.,
see e.g. [10, 11, 12, 13, 14, 15, 16] and, especially, the illuminating summary of those
studies in [3]. Notable exception here were the predictions of
the n as resonant state [17, 18], strongly
contradicting the other studies. The understanding of all these research results
is currently quite controversial and probably it will not be easily settled in
the near future.
As a way out of the too complicated theoretical discussions, a straightforward
idea of how to speculate about n in unequivocally tractable terms of
observables was formulated in [19]: let’s find “prearranged”
n in the “atmosphere” of neutron-rich system (e.g. He) and attempt
to populate n continuum by the knockout of the remnant nuclides
(-core in the case of He). Quite low-energy responses (e.g. as
low as MeV) were predicted in [19] for the specific
peripheral character of the reaction mechanism. Such n “ground states”
cannot be interpreted as real resonances, but still can be observed
experimentally and interpreted theoretically as a few-body continuum response
combining effects of FSI and initial-state He structure (ISS) via the
reaction mechanism. The formation of such a low-energy continuum response is
partly similar to the mechanism of the “soft dipole mode” formation in halo
nuclei: (i) very weak FSI effects and (ii) strong low-energy concentration of
the strength function due to the peripheral character of both the halo wave
functions (WF) and electromagnetic transition operators
[20, 21, 22]. In [23] the
data of [6] were analyzed in analogous assumptions to Ref. [19]. However, conclusion made in [23] that
the He-induced source can directly explain the data of [6]
without any peripheral assumptions, is in a qualitative contradiction with the
results of [19].
In this work we demonstrate that an evidence for the low-energy structures
analogous to the observation of [5, 6] can be found in
the other reactions with He beam. As for the theoretical discussion we
revisit the n population from He in simple approach
[19], but at a modern technical level. We generally confirm
the results of [19], but provide an additional insight into
the reaction mechanism, connected with specific correlations in the He WF.
Possible reasons for disagreement with the results of [23] are
analyzed. The system of units is used in this work.
2 Experiment
Experiment with the He beam impinging the deuterium target was performed
at the ACCULINNA-2 facility (FLNR, JINR), having the search for the H
nuclide as a goal [24, 25]. The experiment appeared
to be much more “treasurous” than intended, providing interesting results also
on H [26] and other auxiliary channels like Li,
Li [27, 28] and H, He, He
[29]. Here we present the results obtained for the
H(He,Li) and H(He,He)HH+
reactions, previously omitted. Since the experiment has been already well
presented in literature, we only briefly describe here the most relevant
details.
Fig. 1: (a) Identification of Li recoil nuclei (red dots) by - method in the side telescopes. (b) The ToF distribution obtained for the
stilbene-array signals triggered by the side telescopes. The set of three peaks
with ToF ns corresponds to the gamma rays produced in the diaphragm
installed 20 cm upstream the target plane, in the target frame, and in CsI(Tl)
array. The green and red histograms are formed by the events, identified by the
-TAC method as gammas and neutrons, correspondingly.
The He beam was produced at 26 A MeV with intensity pps.
The beam was focused on the cryogenic gaseous deuterium target, with a
temperature of 27 K, equipped with the thin stainless-steel and mylar windows.
The target thickness was . The setup allowed us to measure the H(He,Li)n
reaction and some products of the multibody decay of the produced unbound
system.
The beam particles were identified by the two plastic scintillators, allowing
to deduce the energy of the projectile from its measured time-of-flight (ToF).
The projectile trajectories were reconstructed by two pairs of multi-wire
proportional chambers. The special run with the empty target cell was performed
to estimate the background conditions, which had of the total
He beam time.
The recoil Li nuclei, appearing in the H(He,Li) reaction hit the
array of four identical -- telescopes. The telescope array was
located 179 mm downstream the target. Each telescope consisted of three layers
of silicon strip detectors (SSDs). The 20-m-thick SSD with a sensitive area
of mm was divided into 16 strips, the second and the third
layers were created by the two identical 1 mm-thick SSDs ( mm
with 16 strips). The Li recoils emitted from the deuterium gas target in the
H(He,Li) reaction in a range degrees in the lab system were
detected by this telescope array. The telescopes allowed to identify Li with
clear separation from the other registered lithium isotopes, see Fig. 1 (a). The central telescope was installed at the beam line at the
distance of 323 mm behind the target. It was intended to detect tritons emitted
with high energies at the angles smaller than in the laboratory
system. The telescope consisted of one 1.5-mm-thick double-sided SSD ( mm, with 32 strips on each side) followed by a square array of 16
CsI(Tl) crystals. The crystals had a cross section of mm
and thickness 50 mm each.
The group of four neutrons appears in “free flight” as a result of the
-core removal from the He projectile. An important part of the
H(He,Li)n reaction analysis was the neutron identification
and reconstruction. The neutron-wall setup [30] included 48
stilbene scintillator crystals placed on a m area located 2
m downstream the target at zero angle to the beam axis. The distance between the
50-mm thick and 80-mm diameter stilbene crystals was approximately 12 cm which
resulted in about probability for neutrons to hit a stilbene detector.
The stilbene array provided energy resolution and the single neutron
registration efficiency of . The probability of a neutron
registration in coincidence with the Li recoil was around , taking
into account that four neutrons are flying forward, towards the stilbene array.
The -gamma separation was made by means of the Pulse Shape Discrimination
(PSD) [30]. The PSD information, augmented with ToF
distribution, see Fig. 1 (b), leads to suggestion that some
gamma-type signals correspond to neutron-like ToF. These gammas are presumably
produced by neutrons interacting with the stilbene modules housings and should
be taken as neutron events.
2.1 H(He,Li)n results
Tetraneutron was reconstructed from the recoil Li as a missing-mass (MM)
component in the H(He,Li) reaction. The total number of Li-
coincidences found in the recorded data was 136, see Fig. 2 (a).
In this figure the neutron kinetic energy in the n c.m. frame is
compared with the reconstructed MM energy of the 4n
group. The fact that the majority of these events (108 events) are
located inside the kinematically allowed region proves the good channel
identification for the data. The empty-target measurement Fig. 2
(a) shows that the background conditions were very “clean” for the n
population in the H(He,Li) reaction: the “empty” events are very
few and located mainly in the unphysical part of the kinematical plane.
Fig. 2: (a) Correlation between the neutron energy in the n frame and the
n decay energy for the H(He,Li) reaction. The kinematical
limit is shown by
the red dashed line separating the accepted events (red diamonds) and rejected
events (black diamonds). The empty-target measurement events are shown by green
circles. (b) The n MM spectrum. Green dots show the
simulated empty-target spectrum. The blue dashed curve corresponds the 4-body
phase volume and thick gray curve to “Fourier transform” of
the He source.
The MM spectrum of n, Fig. 2 (b), was reconstructed from
the measured Li recoil energy and emission angle, taking the 108 events with
the data points located within the “kinematical triangle” . The obtained spectrum shows a group of 6 events around
MeV. The width of this group is well consistent with the
value of energy resolution MeV in this region of n MM spectrum.
The statistics of this low-energy group is marginal, but resolution is
sufficient and background conditions are shown above to be very good.
Obviously, the observed low-energy hump in Fig. 2 (b) can not be
described just as a contribution of the 4-body phase volume or “Fourier
transform” of the He source, see also Fig. 8 (d). The
“generic” 4-body phase volume behavior is , but for the
n population the Pauli principle requires at least two additional
“excitation quanta” for the lowest-energy configuration or in
terms of the hyperspherical harmonics (HH) model. The hypermomentum is grand
quantum number in the few-body configuration space. This leads to a modified
4-body phase volume behavior . It is notable that the observed
energy of this group events ( MeV) is consistent (within the
experimental energy resolution) with the n low-energy peak value (sys.) MeV reported in Ref. [6].
2.2 H(He,He)HH+ results
Fig. 3: (a) “Kinematical triangle” for n MM vs. H MM for the
H(He,He)H data of [25]. Green dots show the
empty target events. Red dashed and blue dotted frames show the event selection
for Fig. 4 (b). (b) Tetraneutron MM spectrum summed up for
MeV. Large green circles show the simulated empty-target
spectrum. Thick gray line is MC simulated phase volume for H decays.
In Ref. [25] the excitation spectrum of H was
populated in the H(He,He)H reaction up to MeV above the 5-body H+ decay threshold. An evidence was reported
for resonant states at and 5.5(3) MeV and also some
indications for such states at 7.5(3) and 11.0(3) MeV. These structures were
observed at small c.m. reaction angles . The
H excitations in a broad range with
MeV form quite a featureless hump up to maximal available
energy. It is possible to search for the low-energy-correlated n emission
off these sufficiently highly-excited configurations of H.
Fig. 4: (a) “Kinematical triangle” for n MM vs. Li) MM
for the H(He,He)H data of [25]. Large green
circles show the empty target events.
(b) MM spectrum (black solid histogram). Red dashed and
blue dotted histograms show spectra corresponding to selection frames in Fig. 3 (a). (c) Tetraneutron MM spectrum gated on the lowest states
in the He+H continuum of the system, see
frames of corresponding style in the panel (a).
The data of Ref. [25] contain coincidence information on the
H decay product of H, which allows to infer MM of the n
subsystem. These data in terms of the n emission are shown in Fig. 3 (a). The H events are mainly concentrated close to
diagonal , which means that the “internal” energy of
the tetraneutron tends to be large, while the triton gets only a small fraction
of the total H decay energy. This situation seems to be close to the phase
volume distribution of the 5-body H+ decay for given H decay
energy as
in terms of n energy-distribution parameter . The standard
value for 5-body phase volume is . It was discussed in the previous
section that at least value is requested for by the Pauli principle for n. This idea is confirmed by the
calculations of the 5-body decay of H in [31]. The
distributions at given H energies are also quite “compressed” due to
experimental bias, see Fig. 13 of [25], so for
distribution Eq. (1) turns out to be completely justified in our
situation.
To observe the possible n correlation, we need to find the kinematical
region, where the main “phase volume” component of the decay distribution Eq. (1) is well suppressed. This means , and
for the studies of the n excitation region MeV we need
to consider the H decay energy region MeV. Under this
selection condition, the hump with MeV can be found in the
n MM spectrum, see Fig. 3 (b). The Monte Carlo (MC)
simulated phase volume only weakly “penetrates” in the MeV
region. The background of about events/MeV (green dots in Fig. 3 (b) is estimated from the available empty target data. Thus
significant part of this low-energy n data may be connected with
the random background.
Fortunately, it appeared that the low-energy n events in Fig. 3 (b) are nicely correlated with the population of the lowest
MeV state in the He+H continuum of the
[32], see Fig. 4 (b).
This means that we are actually dealing here not with the HH+ decay, but rather with the H(He,Li reaction.
Moreover, the same correlation is perfectly true not only for events in the
MeV range, but also for the MeV range. In
this range most of the events should be either unphysical or connected with low
energy resolution, as the population of the MeV range is
predicted to be negligible in all theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy
resolution for the n MM spectrum is MeV FWHM. However, this
resolution has a broad “nongaussian” component, which may result in some
number of n events with MeV. The coincidence with
definite state in Li practically guarantee the physical nature of such
event, so we recognize that absolute majority (26 events out of 28) found in the
range MeV in Fig. 3 (b) as really
belonging to n MM spectrum.
Strong correlation found in Fig. 4 (b) inspired us to
“reverse” the logic and construct the n spectrum gated by the lowest
states in the continuum, see Fig. 4
(c). There is a strong sensitivity of the obtained n MM spectrum to the
energy of the populated Li states. The low-energy n spectrum is
populated only for MeV. By comparing the dashed
and dotted histograms in Fig. 4 (c), one may find indication
that “high-energy” parts of n spectrum are qualitatively different for
the ranges and MeV. This
may be indication that within these energy ranges there are actually two states
of Li: and MeV [32]. So far this
speculation should be taken with caution because of concentration of “empty
target” events in the high n MM energy part of these spectra, see Fig. 3 (a).
3 On n population in different reactions
The low-energy population in [6] and in the processes considered in
this work all have the He nuclide as a starting point. At first glance,
this is the only similarity, and the reaction mechanisms should be too
different. Let’s demonstrate that the differences are actually not so striking.
Table 1 shows the momenta and velocities in the recoil-product
channel for the reaction of Ref. [6] and reactions considered in
this work. Of course, in the reaction of [6] the “recoil”
particles are leaving the FSI region faster, than in the case of our Li)
reaction, but in the timescale this “faster” is only a factor of 2.
Fig. 5: Momentum transfer to n system for the H(He,Li)n
reaction are shown by black circles in (a). The same for the
H(He,He)HH+ reaction: blue diamonds and
red stars in (a) corresponds to ranges
and MeV, blue diamonds and red stars in (b) corresponds to
and MeV. Large green circles show the corresponding
empty target events.
Table 1: Relative momentum and relative velocities in the
recoil-product channel for reactions of interest.
The H(He,He)HH+ process formally has the
decay process as the second step. In fact, from Table 1 it is
clear that for the excited states of H, which we consider, the velocity of
the H recoil leaving n in the “decay process” is actually
comparable with the velocity, at which the He recoil had left the FSI region
in the first [the H(He,He)H reaction] step of this process. This
kinematical region is actually a transition region from
H(He,He)H reaction to H(He,n,
where highly excited states of Li (above the He-H threshold)
are populated. We find in Fig. 4 (c) that the most expressed
low-energy n indication is obtained for the recoil,
which is the lowest He-H resonant state.
So, we see that the low-energy n continuum is confidently populated in the
H(He,n reactions with excitation energies
MeV, see Fig. 2 (b), and
MeV, see Fig. 4 (c). It is
quite natural to expect that such a population should also take place at the
intermediate energies, which correspond to + continuum states of the
Li. These would be reactions of quasifree scattering H(He,)n, which have never considered seriously because of extremely
“soft” H projectile. So far, the possibility to extract this information
from our existing data is questionable.
Another characteristic, which is expectedly very important for the direct
reactions, is momentum transfer . For our reactions the
values are easily reconstructed, see Fig. 5. The momentum
transfers MeV are realized, what is found important for
data interpretation, see Fig. 7 and related text.
Unfortunately, we do not find information on the momentum transfers in Ref. [6], though this information is obviously accessible in these data.
4 Theoretical model
Table 2: Geometry of He (rms radial characteristics in fm) versus
experimental data.
Th.
Exp.
In this work we performed studies of He in a five-body + and
of n in a four-body hyperspherical harmonics models. The detailed
account on these studies will be given elsewhere, and here we focus on the
aspects relevant to n production from He, following the ideology of
Ref. [19]. The HH Schrödinger equation (SE) formulation
for particles is used either without (bound states) or with right-hand-side
inhomogenity (continuum states):
On the properly antisymmetrized HH basis
the variational procedure reduces the SE with inhomogeneous term
(2) to a set of coupled ordinary differential equations:
where is “multiindex” numbering the available HH basis states
and
For the SE (3) without Coulomb interaction the boundary
conditions at large are known analytically
where is hypermomentum, are the modified
Bessel
functions of the second kind, are the Riccati-Bessel functions.
The “source functions” are terms of
hyperspherical expansion of the inhomogeneous term :
The easiest reaction model which can be considered for n production is
“sudden removal” of core from He [19, 23],
defined by action of annihilation operator on the -core. This procedure
defines the “source function” in (2) as the
Fourier transform of the overlap integral between the -cluster WF
and the He WF over the radius-vector
between the removed -cluster and the center-of-mass of n:
The “strength function” for population of the n continuum should be
proportional to the outgoing flux of particles, expressed via the WFs with
pure outgoing wave boundary conditions Eq. (5)
found on a hypersphere of a large radius fm.
5 Properties of the He
Following the cluster ++ model studies of He, Li,
Be isobar [33, 34, 22] we use the
Sack-Biedenharn-Breit (SBB) potential [35] in the -
channel. We suppress the Pauli forbidden state by using the additional repulsive
core in -wave potential, which provides practically the same scattering
phases as the deep potential.
Table 3: Elastic and reaction cross sections for He (in mb).
In the calculations we adopted the local charge independent Argonne SSC AV14
potential [38]. For evaluations we also used simple -wave
singlet potential Brown-Jackson (BJ) from [39]: , MeV, fm. For three nucleons with separations
we use the three-nucleon potential [40]
in Eq. (4). A set of the parameters MeV, , MeV, yields the
binding energies 8.41 (8.48), 7.74 (7.72), and 28.44 (28.30) MeV for H,
He, and He ground states, respectively (the experimental values are
shown in parentheses).
Fig. 6: (a) Correlation density of the He WF. (b) Correlation density of
the n source component with with qualitative
illustrations of dominant correlation patterns. (c) Momentum-space correlations
for dominant , component of the source (b).
The calculations were performed with , containing 15862 basis
states reduced to 675 states in (3) by antisymmetrization.
The percentages of the major He WF components are
are in a good agreement with advanced ab initio Quantum Monte Carlo
calculations of [41] and [42]
. The geometric characteristics of He are compared with
experimental data in the Table 2. The root mean square (rms) charge
radius of He, has been determined for the first time as
1.93(3) fm [43]. However, using the new charge radius
1.67824(83) fm of He [44] as an anchor point for the isotope
shift measurements, the He charge radius was reevaluated as 1.9559(7)(158)
fm, where the first uncertainty is from new charge radius value and the second
uncertainty from the electronic isotope shift measurements. Our result for
He charge radius is in nice agreement with both experiments. The comparison
with experimental in Table 2, seems to be not so
favorable, but we need to recall that this value is actually extracted from the
high-energy reaction data in a model-dependent way. Therefore, here the direct
comparison of cross sections is more preferable. The results of Glauber-like
calculations with He densities obtained in our model are shown in the
Table 3 to be nicely consistent with experimental data, including
the cross section on the C target [37], from which the
matter radius of fm, shown in Table 2 was actually
extracted.
6 The He WF as a source for n population
The He WF correlation density vs. (informative for
the n studies) is shown in Fig. 6 (a). It looks
overall quite featureless. In contrast, the projection of the He WF on the
n configuration with , where strongest n FSI
effect is expected, exhibit strong spatial correlations, see Fig. 6 (b). These correlations can be interpreted basing on a
simple relation between rms radius of individual neutron in the n c.m. frame and rms hyperradius
Following the well-known naming of spatial correlations in the classical
He halo nucleus — “dineutron” and “cigar” [33] — we
characterize correlations in He source as “broom” (compact tetraneutron
aside of -particle) and “cross” (neutrons evenly distributed around
central-located -particle). The double-hump picture of correlations can
be qualitatively
understood as connected with configuration (in the shell model
notations) in analogy with configuration providing strong spatial
focusing (“Pauli focusing” [33]) in the He case. Fig. 6 (c) shows the Fourier transform of the source (b),
illustrating the momentum content of this formation without n FSI.
It should be understood that the strongly correlated source shown in Fig. 6 (b) is related to about in the He WF. The
other components of these WF projected on n system populate “excited”
configurations of n, where we do not expect any strong final-state
effect.
Fig. 7: (a) The n spectra obtained for MeV without FSI,
with reduced - FSI ( fm), full n FSI with normal source
and with peripheral source (black solid curve). (b) n spectra obtained for
different transferred momenta for peripheral source.
Figure 7 (a) shows the evolution of n strength
functions for selected MeV: no FSI reduced FSI
full n FSI. The rms size of the source fm presume significantly smaller distances,
than in initial He WF. The strength function peak for full n FSI is
at about 14 MeV. To form extreme low-energy peak in the strength function in the
region MeV as indicated by data, a further
qualitative assumption of the peripheral character of reaction is needed. To
evaluate it, we cut out small hyperradii of the original source function by a
profile function of the form
and further refer it as “peripheral source”. With fm and
fm the rms size of the source is increased to fm and the desired energy position of the peak is obtained.
The systematics of the strength functions for peripheral sources for different
momentum transfers is illustrated in Fig. 7 (b). One may see that even the “peripheral assumption” is
not sufficient to form a strong and distinct low-energy peak in the n
spectrum: the desired picture is realized only for certain intermediate range
of transferred momenta MeV. This range well corresponds
to situation of our experiment, see Fig. 5.
A note about convergence. The calculations are performed up to ,
which is insufficient for complete energy convergence. For that reason we
perform simultaneously the calculations with advanced AV14 potential and simple
central - potential BJ with much faster and “safer” convergence. The
convergence curves are used for exponential extrapolation to infinite basis. For
example, for calculations of Fig. 7 (a), the convergence
point was MeV, while “underbinding” was MeV for AV14
and MeV for BJ potentials. For Fig. 9
the corresponding values are , and MeV.
The additional control of the energy may be provided by phenomenological
few-body potential with Fermi-function profile
The quite “soft” 4-body potential (11) with parameters
fm and fm was used to get the peaks at
energies pointed by the exponential extrapolation. “Safety” of the procedure
in the sense of strength function shape was checked by
extrapolations.
Qualitative comparison of theory and available data is provided in Fig. 8. Here we need to recall that strongly correlated picture of
Fig. 7 is connected with of the He WF,
while of the sources populate excited configurations of the n
are evaluated to provide strength function with maximum at MeV, see
thick gray curve in Fig. 8 (d).
Fig. 8: Qualitative comparison of theory and data. (a) Data Duer et
al. [6], (b) H(He,Li) reaction, (c)
H(He,He)HH+ reaction. (d) Theoretical
results for MeV: nonresonant with peripheral source, resonant
for normal source.
It is also important to note that the low-energy peak connected with extreme
peripheral source has very different profile compared to the case of the real
resonant behavior connected with normal-size source. The “resonant” (red
dashed) curve in Fig. 8 (d) was generated by additional
artificial binding introduced in Hamiltonian by phenomenological 4-body
potential (11) with parameters fm and
fm. The difference between “peripheral” black and “resonant” red curves in
Fig. 8 (d) is too small to be seriously discussed for the
available quality of the data. However, in principle, we have guidelines from
theory how to distinguish experimentally the “peripheral”
reaction-mechanism-induced and “real resonant” tetraneutrons.
7 Theoretical discussion
In the recent paper [23] the data of [6] was
qualitatively reproduced by emission off He-induced source without any
peripheral assumptions. This is in a strong contradiction to the results of
[19] and this work. Let’s try to track the origin of this
contradiction. In Fig. 3 of Ref. [23] a calculation with
reduced - interaction (scattering length fm) is shown.
This calculation is reasonably close to “no FSI” situation and thus easy to
“reproduce”. It is shown in Fig. 9 that a very close result is
obtained in our calculations with simple analytical source
in the , component. The rms neutron radius fm of such a source is very consistent with the values found in
Table 3 of Ref. [23]. The strength function peak shape is
somewhat different from that in Fig. 3 of [23], which
indicates more complicated correlated nature of the source function in Ref. [23], then we assume in Eq. (12). However, the
peak value for the full-calculation n strength function in Fig. 9 well coincide with 2.5 MeV of [23], which
means that we are reasonably “on a right track”.
Fig. 9: Imitation of calculations Lazauskas et al.
[23]. Green dotted curve shows calculation with reduced -
interaction from [23].
So, in principle, there exist a class of source functions which can produce
extreme low-energy peak in the n spectrum for quite compact ( fm) source. The only problem is that the “supersoft” source
Eq. (12) has short-distance asymptotic . If we
get back to the definition Eq. (6), this implies that He
overlap with 4-body hypersphericals is singular as at zero,
which is highly unlikely for any realistic He structure. The n
sources deduced in our work has asymptotic at zero for .
This is in full agreement with analytical results of [19].
Consequently, such sources are much “harder” in the momentum space, and
additional peripheral assumption is needed to get sizable low-energy
effect in n spectrum.
The origin of the “supersoft” source in [23] is probably due
to the fact that n c.m. is assumed to coincide with He c.m. in
Ref. [23]. This is an implicit assumption of infinitely heavy
-core in He and, of course, quite a poor approximation considering
the actual mass ratio of and n. This assumption also leads to
total absence of such an important characteristic as in the
formalism of [23]. In contrast, the “geometric” effects of
n motion relative to in He WF are found to be very
pronounced in our calculations, see Fig. 6.
So, we insist that peripheral assumption is necessary to describe the
extreme low-energy peak in the spectrum of n. This point is also
qualitatively supported by extreme small cross sections for the observed
peculiarities in the n spectra. Typical cross section values for analogous
direct reactions are some millibarns and even tens of millibarns,
while we and Ref. [6] observe some microbarns. Such a cross
section suppression looks quite natural if only the extreme periphery of the
original n configuration is actually participating in the reactions of
interest without destroying rescattering effects.
8 Conclusion
Evidence for a hump in the continuum at about 3.5 MeV was observed in
the two reactions: H(He,Li) and
H(He,He)HH+. We demonstrate that the
reaction mechanism in both cases is actually best interpreted as transfer
reaction to the ground and to the MeV states of Li. The reaction
mechanism also can be found reasonably consistent with that in the
H(He, He) reaction [6], so it could be the
same type of phenomenon. The statistics of 6 events for the first reaction is
quite low, but events for the second reaction is comparable to
statistics connected with the low-energy structure observed in
[6].
Paper [23] is explaining the low-energy tetraneutron peak in
Ref. [6] as solely an initial state (He structure) effect in
the presence of n FSI. This is in a strong contradiction to the results of
[19] and of this work. As compared to [19]
(i) the calculations of this work are performed with realistic NN and 3N
interactions, and (ii) the obtained realistic He source demonstrates the
complicated correlated behavior that is very important for the n
population. However, these two aspects are not sufficient to fully explain the
presence of a low-lying peak in the tetraneutron spectrum. To do this, it is
also necessary to consider extremely peripheral reaction mechanism, in which the
-core of He interacts with the target, but four neutrons remain
“spectators” at distances, far exceeding the typical sizes of neutron orbitals
in He. Using the appropriate profile function of the type
(10), simulating such a geometry, it is possible to explain the
presence of a low-lying peak in the tetraneutron spectrum. Since the appearance
of the low-energy peak is not related to the tetraneutron per se, but
also to the source used and the reaction mechanism, it should be expected that
it could be detected at somewhat different energies in different reactions. A
possible reason of disagreement with the calculations [23] may
be the incorrect n c.m. treatment in He WF constructed in
[23].
The existence of low-energy tetraneutron resonance would mean a radical
revision of everything we know about neutron-rich nuclei and neutron matter. Our
vision of the problem is that a solution can be found, which is much less
radical being related to the He structure and reaction mechanisms. We got
an important indication in our data that the n spectrum is strongly
dependent on the particular state of the recoil Li system, which means,
first of all, on the reaction mechanism. If our understanding of the
low-energy structures in tetraneutron is true, then it is not disappointing at
all not to find tetraneutron resonance, as actually we are at a new
exciting field of studies of extreme-peripheral phenomena in nuclear systems.
It is clear that the further studies of He induced reactions on the
deuteron target are needed to bring to full reliability the evidences obtained
in this work. However, the processes of the major interest are all in the cross
sections range of microbarns. Therefore, such studies make sense only if the
statistical limitations of the existing data are overcome by about an order of
the magnitude, which is a real challenge for available experimental techniques
and facilities.
Acknowledgments
— This work was supported in part by the Russian Science Foundation (grant
No. 22-12-00054). The research was supported in part in the framework of
scientific program of the Russian National Center for Physics and Mathematics,
topic number 6 “Nuclear and radiation physics” (2023-2025 stage). The authors
are grateful to Prof. Yu.Ts. Oganessian for the long-term support of this
research.